OpenStudy (anonymous):

What is the solution to the equation 6y –2(y + 1) = 3(y – 2) + 6?

OpenStudy (anonymous):

y=2

OpenStudy (ilovecake):

6y-2(y+1)=3(y-2)+6 Simple and best practice solution for 6y-2(y+1)=3(y-2)+6 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 6y-2(y+1)=3(y-2)+6 equation: Simplifying 6y + -2(y + 1) = 3(y + -2) + 6 Reorder the terms: 6y + -2(1 + y) = 3(y + -2) + 6 6y + (1 * -2 + y * -2) = 3(y + -2) + 6 6y + (-2 + -2y) = 3(y + -2) + 6 Reorder the terms: -2 + 6y + -2y = 3(y + -2) + 6 Combine like terms: 6y + -2y = 4y -2 + 4y = 3(y + -2) + 6 Reorder the terms: -2 + 4y = 3(-2 + y) + 6 -2 + 4y = (-2 * 3 + y * 3) + 6 -2 + 4y = (-6 + 3y) + 6 Reorder the terms: -2 + 4y = -6 + 6 + 3y Combine like terms: -6 + 6 = 0 -2 + 4y = 0 + 3y -2 + 4y = 3y Solving -2 + 4y = 3y Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '-3y' to each side of the equation. -2 + 4y + -3y = 3y + -3y Combine like terms: 4y + -3y = 1y -2 + 1y = 3y + -3y Combine like terms: 3y + -3y = 0 -2 + 1y = 0 Add '2' to each side of the equation. -2 + 2 + 1y = 0 + 2 Combine like terms: -2 + 2 = 0 0 + 1y = 0 + 2 1y = 0 + 2 Combine like terms: 0 + 2 = 2 1y = 2 Divide each side by '1'. y = 2 Simplifying y = 2

OpenStudy (ilovecake):

y = 2

OpenStudy (anonymous):

preciate it